A treasure map is drawn to a scale of 2 inches = 3 miles. On the map, the distance between Point A and X-marks-the-spot is 9 1/2 inches. How many actual miles does this represent.
Answer by mananth(16946)

2 inches = 3miles
9 1/2 inches = 19/2 inches = ?
..
19/2 *3 /2
= 57/4 miles
=14.25 miles